# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ Abstract.
We study an individual-based stochastic SIR epidemic model with infection-age dependent infectivity on a large random graph, capturing individual heterogeneity and non-homogeneous connectivity. Each individual is associated with particular characteristics (for example, spatial location and age structure), which may not be i.i.d., and represented by a particular node. The connectivities among the individuals are given by a non-homogeneous random graph, whose connecting probabilities may depend on the individual characteristics of the edge. Each individual is associated with a random varying infectivity function, which is also associated with the individual characteristics. We use measure-valued processes to describe the epidemic evolution dynamics, tracking the infection age of all individuals, and their associated characteristics. We consider the epidemic dynamics as the population size grows to infinity under a specific scaling of the connectivity graph related to the convergence to a graphon. In the limit, we obtain a system of measure-valued equations, which can be also represented as a PDE model on graphon, which reflects the heterogeneities in individual characteristics and social connectivity.
@@ 1. Introduction
In the current paper, we are interested in the large population limit of a stochastic model of an epidemic that spreads among an heterogeneous population with random connectivities. Such models have been studied to an extent in the literature. For example, in [11], an epidemic model with an age structure and various social activity levels is studied to understand the effect of population heterogeneity on herd immunity. In addition to ‘age’ as an obvious heterogeneous factor, spatial locations and other characteristics/features may also indicate individual heterogeneity. Another source of heterogeneity arises from the individual connectivities (such as households, communities and social activities), which is often modeled as a non-homogeneous random graph. For example, epidemic models on random graphs with given degrees, typically under the configuration model are studied in [1, 53, 50, 9, 5, 33, 25, 40, 42, 15], on weighted (configuration) graphs [12, 18, 13, 51], on dynamic (evolving) graphs [3, 2, 34, 22, 14, 24, 29]. (See also the relevant models on random networks with household structures in [7, 4, 6, 39], and on multilayer networks in [30, 39].)
@@
A few papers have established large population limits for epidemic models on random graphs. In particular, [17, 32] proved a functional law of large numbers (FLLN) for a Markovian SIR model on a configuration model graph with specified degree distributions and edges being randomly matched, and established the measure-valued limit as a systems of nonlinear differential equations, which verifies the conjecture in [53]. In [38], a functional central limit theorem (FCLT) is established for a similar Markovian SI model for the total count processes. In [10], an FLLN is established for a Markovian SIR model on a stochastic block model. In [17, 32, 38, 10], the degree distribution converges to a finite limit as the number of nodes tends to infinity and the number of edges is thus of the same order as the number $n$ of nodes. Such a description with very sparse graphs leads to limits that are of distinct nature from the one of dynamics on a graphon. On the other hand, for dense connectivities among individuals, one can derive large population approximations of the epidemic dynamics as PDEs on graphons. In [37], an FLLN is established for density-dependent Markov processes with finite state space on large random graphs, which includes the Markovian SIS model with individual heterogeneity on graphs as a special case, and a PDE on graphon is derived as the limit. In [20], an FLLN is established for a Markovian SIS model with a general form of individual heterogeneity on random graphs, where a PDE on graphon for the measure-valued state descriptors is also derived. In the two later references ([37, 20]) the scalings of the parameters allow for the random generation of graphs whose number of edges is of order $n^{1+a}$, where $n$ is the number of nodes and $a$ can be freely taken in $(0,1]$. This extends the classical assumption of a dense graph for the convergence to a graphon, where the number of edges is of order $n^{2}$. However, all these FLLN results are about epidemic models on random graphs that are Markovian, where the infectious durations or recovery times are exponentially distributed. In the present paper, we establish an FLLN for a non-Markovian SIR model on large random graphs that results in a measure-valued limit on a graphon.
@@
Non-Markovian stochastic epidemic models with a homogeneous population (no age structure, homogeneous social connectivity, etc.) have been recently studied, see the recent survey [28]. Although these models offer much more possibilities to fit observational data on infection profiles, many of the classical tools in probability cannot be directly exploited. In particular, the standard epidemic models with a constant infection rate and a general infection duration distribution are recently studied in [44] (see also [55, 54] for Gaussian approximations and [49] for measure-valued state descriptors and PDE limit for the FLLN), and epidemic models with varying infectivity are studied in [26, 45, 27], where deterministic or stochastic integral equations are derived for the FLLNs and FCLTs, respectively. By tracking the elapsed infectious times, measure-valued processes are used to describe the epidemic evolution dynamics in models with such infection-age dependent infectivity, and the corresponding PDE and SPDE limits are established for the FLLNs and FCLTs in [46] and [47], respectively. An epidemic model with contact tracing and general infection duration distributions is studied in [21], where measure-valued processes are used to describe the dynamics and the FLLN is established with a PDE limit. In [39], an individual-based multi-layer SIR model with households and workplaces to take into account social connectivity heterogeneity is studied, where measure-valued processes tracking remaining infectious times are used to describe the epidemic dynamics and a PDE limit is established for the FLLN.
@@
In this paper we consider an SIR model with infection-age dependent infectivity on a large random graph that captures both individual and connectivity heterogeneities. In particular, connectivities among individuals are given by a non-homogeneous random graph. Each node on this graph represents a single individual, each associated with an individual characteristic/feature. These may account for spatial locations, age and/or social belongings. So they may be distributed on a compact subset of ${\mathbb{R}}^{d}$, as in the spatial SIR model considered in [36], which we generalize in the simplest setting of pairwise interactions. In the case of age structure or social activity levels, our model generalizes [11] in that they only consider a finite number of compartments while our model includes a possibly continuous age model. The random connectivity of each edge is then assumed to depend on the individuals’ characteristics in a general manner. The connectivities indicate various levels of social activities, which may for instance depend on the spatial locations or age (again more generally than the model with age and social activity heterogeneities in [11]). The random graphs that we consider are of the same form as the ones in [20]. The main difference is that the connectivity of each edge was assumed in [20] to be a deterministic function of the pair of individuals’ characteristics, while we allow for an additional degree of randomness. In [37], there was no heterogeneity in the contact rate and the dependency of the graph on individual types is more restrictive. Our model extends the framework of stochastic block models considered in [10] except that the scalings of the contact probability are compatible only in the degenerate cases.
@@
Each individual is moreover associated with a random varying infectivity function, which reflects the propensity for/ force of infection of each individual at any elapsed time of infection, and depends as well on the individual characteristics. As a result, the infectious duration is also individualized (we express the corresponding distribution as a function $F_{x}(\cdot)$ of the individual characteristic $x$).
@@
Individuals are grouped into three compartments: Susceptible, Infected and Recovered. Infections are generated by the interactions between susceptible and infected individuals. Each individual is thus exposed to a specific force of infection which changes over time and which is given by aggregating the weighted infectivity levels of the individuals that are connected to him/her in the graph. The whole dynamic of the epidemic depends on the non-homogeneous random graph and on the random infectivity functions through the expression of this total force of infection for each individual.
@@
To describe the evolution dynamics, we use three measure-valued processes for the three compartments, where the measure for the infected process is over both the individual characteristics and the infection ages, while the measure for the susceptible and recovered processes is only over the individual characteristics. We prove a FLLN for these measure-valued processes when the population size goes to infinity. As the population size increases, so does the connectivity graph. We impose conditions on the connectivity probabilities to keep the graph consistency as it grows, which notably covers the case of a graphon in the limit. The limits for the FLLN are given by a set of measure-valued equations, from which we further derive a PDE model for the measure-valued infection process. The PDE model is linear, but with a boundary condition which is given by the product of the measure-valued susceptible process and the aggregate force of infection. To note, the dynamics described from the set of measure-valued equations and the PDE may be seen as evolving on a graphon.
@@
The proof for the FLLN is challenging with this level of realism because of the complicated dependence among the variables and processes. First of all, the individual characteristics are not assumed to be i.i.d., and they affect the random infectivity functions and the edge connectivities. Second, the measure-valued processes all depend on the aggregate forces of infection that act specifically on each individual. Notably, these forces of infection are not functions of the three measure-valued state descriptors, but more generally expressed in terms of the detailed graph structure. Thus it is difficult to prove the convergence of the processes directly.
@@
Our approach consists in constructing an intermediate model for which the infection rates are derived from the limiting model of the FLLN rather than the interactions through the connectivity graph. This approach facilitates the proof of convergence by enabling specific decoupling of the dependence. It extends the one that was first developed for the homogeneous model in [27], which was motivated by ideas from the theory of propagation of chaos [52]. In this way, the regularity conditions imposed on the random varying infectivity functions in [26] could be relaxed to only require their boundedness. The current extension is highly non-trivial however because of the heterogeneities among both the individuals and the connectivities. Moreover, we consider measure-valued processes in this paper, while in [27], only total count processes are considered. Different weak convergence criteria are therefore required for this setting.
@@
In addition, we have made the effort to allow for the individual characteristics to take values in a state space as general as possible instead of the classical choice of $[0,1]$ for graphon kernels. This is particularly of importance regarding the regularity assumptions of the interaction kernel in the limit, which we only assume to be almost everywhere continuous. One of our motivations was to allow for kernels on multidimensional state spaces that possibly display discontinuity thresholds in terms of the distance between the two parameters, as discussed in [20]. We prove in Appendix A a generic convergence result of independent interest, to deal with these weak regularity properties in conjunction with weak convergence of random measures in probability. The approach demonstrates its potential by extending very efficiently to this level of heterogeneity, even with the randomness additionally considered for the varying infectivity functions. The conditions that we assume on the average infection rate per edge and on the average variance in this infection rate seem very close to optimal. We refer the reader to Section 5 for the detailed construction of the intermediate processes and for the proofs of weak convergence.
@@ 1.1. Organization of the paper
The rest of the paper is organized as follows. We give a summary of some common notations used throughout the paper in the next subsection. We provide a detailed model description and assumptions in Section 2. In Section 3, we present the main result of the paper. We establish the uniqueness of the solution to the limiting deterministic PDE model and establish some properties of that solution in Section 4. The proofs for the convergence of the FLLN are given in Section 5, with additional technical supporting results proved in Appendix A.
@@ 1.2. Notation
For a measurable set $(\mathcal{X},\mathcal{F})$, we denote by $\mathcal{M}_{1}(\mathcal{X})$ and $\mathcal{M}(\mathcal{X})$ the sets of respectively probability and non-negative finite measures on $\mathcal{X}$. For $\mu\in\mathcal{M}(\mathcal{X})$ and a real-valued measurable function $f$ defined on $\mathcal{X}$, we will sometimes denote the integral of $f$ with respect to the measure $\mu$, if well-defined, by $\langle\mu,\,f\rangle=\int_{\mathcal{X}}f(x)\,\mu(\mathrm{d}x)=\int f\,\mathrm{d}\mu$. For a metric space $\mathcal{X}$ endowed with its Borel $\sigma$-field, we endow $\mathcal{M}_{1}(\mathcal{X})$ with the topology of weak convergence. We also denote by $\mathcal{D}(\mathbb{R}_{+},\mathcal{X})$ the space of right-continuous left-limited (càd-làg) paths from $\mathbb{R}_{+}$ to $\mathcal{X}$. This space is endowed with the Skorokhod topology (see, e.g., [8, Chapter 3]).
In what follows $(\mathbf{\mathbb{X}},\mathrm{d})$ will denote a Polish metric space (complete and separable).
@@ 2. Model description
We consider a population consisting of $N$ individuals, who are in either of the three states – susceptible, infected or recovered at each time, and may interact with each other according to a random graph described below. Susceptible individuals can be infected randomly through interactions with the infected ones. Once infected, an individual will become infectious for a random period of time until recovery, and once recovered, he or she will no longer infect any susceptibles, or become infected again. Let $\mathcal{S}^{N}(t),\mathcal{I}^{N}(t),\mathcal{R}^{N}(t)$ be the subsets of $[\![1,N]\!]$ that denote the sets of susceptible, infected or recovered individuals at each time $t$, respectively. The corresponding processes $S^{N}(t)=|\mathcal{S}^{N}(t)|$, $I^{N}(t)=|\mathcal{I}^{N}(t)|$ and $R^{N}(t)=|\mathcal{R}^{N}(t)|$ denote the numbers of susceptible, infected or recovered individuals at each time $t$.
@@ 2.1. Infection-age dependent infectivity
For an individual $i\in\mathcal{S}^{N}(0)$, let $\tau_{i}^{N}>0$ be the time at which he/she starts being infected, and let $A_{i}^{N}(t)=t-\tau_{i}^{N}$ be the infection age of the individual $i$ at time $t$ (by default it is equal to zero for $t<\tau_{i}^{N}$). For an initially infected individual $j\in\mathcal{I}^{N}(0)$, let $\tau_{j}^{N}=-A_{j}^{N}(0)$ denote the infection time before time 00, so that $A_{j}^{N}(t)=t+A_{j}^{N}(0)$ is the infection age at time $t$.
@@
At time $t$, each individual $i\in\mathcal{I}^{N}(t)$ who is infectious has an infectivity function depending randomly on the infection age:
$\lambda^{N}_{i}(A_{i}^{N}(t)),\quad t\geq 0.$ (2.1)
For the initially infected individuals $j\in\mathcal{I}^{N}(0)$, we also define their infectivity level at a given time $t>0$ as $\lambda^{N,0}_{j}(t)$:
@@
A typical way to define these random functions $\lambda^{N}_{i}(\cdot)$ is specified as follows, although we do not rely in our proofs on this representation. Let $\mathbf{\mathbb{Y}}$ be a Polish space. Let $\mathcal{Y}^{N}=\{Y_{i}^{N}\}_{i\leq N}$ be an i.i.d. sequence of random variables taking values in $\mathbf{\mathbb{Y}}$, with common distribution $\mu_{Y}$. We assume that the sequence $\mathcal{Y}^{N}=\{Y_{i}^{N}\}_{i\leq N}$ is independent from $\mathcal{F}_{0}^{N}$. Let $\widehat{\lambda}$ be a deterministic measurable function from $\mathbf{\mathbb{X}}\times\mathbf{\mathbb{Y}}\times\mathbb{R}_{+}$ to $\mathbb{R}_{+}$. Then we can define the random functions $\lambda^{N}_{i}(\cdot)$ for each $i$ as
@@
$\eta^{N,0}_{j}:=\sup\{a>0:\lambda_{j}^{N}(A^{N}(0)+a)>0\}\,.$
We specify the individual distribution of infection duration through its cumulative distribution. Under the formalism of Remark 2.1 it is to be thought as depending only on the variables $(X_{i}^{N},Y^{N}_{i})$.
@@
There exists a measurable function $(F_{x}(a))_{x\in\mathbf{\mathbb{X}},a\in\mathbb{R}_{+}}$ with values in $[0,1]$ which is right-continuous, non-decreasing in the variable ‘$a$’ such that $\lim_{a\to\infty}F_{x}(a)=1$ for all $x\in\mathbf{\mathbb{X}}$ and that the two following identities hold for any $N\geq 1$, $t>0$, $i\in\mathcal{S}^{N}(0)$ and $j\in\mathcal{I}^{N}(0)$ with the notation $F^{c}_{x}(a)=1-F_{x}(a)$:
@@
Recall that $\mathbb{P}_{0}^{N}(\eta^{N}_{i}>t)$ equals by definition $\mathbb{P}\big{(}\eta^{N}_{i}>t\,\big{|}\,\mathcal{F}_{0}^{N}\big{)}$ and similarly that $\mathbb{P}_{0}^{N}(\eta^{N,0}_{j}>t)=\mathbb{P}\big{(}\eta^{N,0}_{j}>t\,\big{|}\,\mathcal{F}_{0}^{N}\big{)}$.
@@
This assumption appear more explicitly stated under the formalism presented in Remark 2.1. By definition, there exists a deterministic measurable function $g:\mathbf{\mathbb{X}}\times\mathbf{\mathbb{Y}}\to\mathbb{R}_{+}$ such that $\eta^{N}_{i}=g(X^{N}_{i},Y^{N}_{i})$. For each $x\in\mathbf{\mathbb{X}}$, we define the cumulative distribution:
@@
For any $N$ and any initially infected individual $j\in\mathcal{I}^{N}(0)$, the constraint that $\eta^{N,0}_{j}>0$, which translates into $g(X_{j}^{N},Y^{N}_{j})>A_{j}^{N}(0)$, explains the form of the second identity in Assumption 2.4, where for any $t>0$,
@@
There exists a function $\bar{\lambda}$ in $\mathcal{D}(\mathbf{\mathbb{X}}\times\mathbb{R}_{+},\mathbb{R}_{+})$ such that the two following identities hold for any $N\geq 1$, any $a\geq 0$, any $t\geq 0$, any $i\in\mathcal{S}^{N}(0)$ and any $j\in\mathcal{I}^{N}(0)$:

### v2

# Stochastic heterogeneous SIR model with infection-age dependent infectivity on large random graphs

@@ Abstract.
We study an individual-based stochastic SIR epidemic model with infection-age dependent infectivity on a large random graph, capturing individual heterogeneity and non-homogeneous connectivity. Each individual is associated with particular characteristics (for example, spatial location and age structure), which may not be i.i.d., and is represented by a particular node. The connectivities among the individuals are given by a non-homogeneous random graph, whose connecting probabilities may depend on the individual characteristics of the edge. To each individual is associated a random infectivity function of its infection age, which is allowed to depend upon the individual characteristics. We use measure-valued processes to describe the epidemic evolution dynamics, tracking the infection age of all individuals, and their associated characteristics. We consider the epidemic dynamics as the population size grows to infinity under a specific scaling of the connectivity graph related to the convergence to a graphon. In the limit, we obtain a system of measure-valued equations, which can be also represented as a PDE model on graphon, and reflects the heterogeneities in individual characteristics and social connectivity.
@@ 1. Introduction
In the present paper, we are interested in the large population limit of the stochastic model of an epidemic that spreads among an heterogeneous population with random connectivities. Such models have been studied to an extent in the literature. For example, in [11], an epidemic model with an age structure and various social activity levels is studied to understand the effect of population heterogeneity on herd immunity. In addition to ‘age’ as an obvious heterogeneous factor, spatial locations and other characteristics/features may also contribute to the individual heterogeneity. Another source of heterogeneity arises from the individual connectivities (such as households, workplaces, communities and social activities), which is often modeled as a non-homogeneous random graph. For example, epidemic models on random graphs with given degrees, typically under the configuration model are studied in [1, 51, 48, 9, 6, 30, 23, 38, 39, 15], on weighted (configuration) graphs [12, 17, 13, 49], on dynamic (evolving) graphs [2, 3, 32, 21, 14, 22, 27]. (See also the relevant models on random networks with household structures in [4, 5, 7, 37], and on multilayer networks in [28, 37].)
@@
A few papers have established large population limits for epidemic models on random graphs. In particular, [16, 31] proved a functional law of large numbers (FLLN) for a Markovian SIR model on a configuration model graph with specified degree distributions and edges being randomly matched, and established the measure-valued limit as a system of nonlinear differential equations, which verifies the conjecture in [51]. In [36], a functional central limit theorem (FCLT) is established for a similar Markovian SI model for the total count processes. In [10], an FLLN is established for a Markovian SIR model on a stochastic block model. In [16, 31, 36, 10], the degree distribution converges to a finite limit as the number of nodes tends to infinity and the number of edges is thus of the same order as the number $n$ of nodes. Such a description with very sparse graphs leads to limits that are of distinct nature from the one of dynamics on a graphon. On the other hand, for dense connectivities among individuals, one can derive large population approximations of the epidemic dynamics as PDEs on graphons. In [35], an FLLN is established for density-dependent Markov processes with finite state space on large random graphs, which includes the Markovian SIS model with individual heterogeneity on graphs as a special case, and a PDE on graphon is derived as the limit. In [18], an FLLN is established for a Markovian SIS model with a general form of individual heterogeneity on random graphs, where a PDE on graphon for the measure-valued state descriptors is also derived. In the two later references ([35, 18]) the scalings of the parameters allow for the random generation of graphs whose number of edges is of order $n^{1+a}$, where $n$ is the number of nodes and $a$ can be freely taken in $(0,1]$. This extends the classical assumption of a dense graph for the convergence to a graphon, where the number of edges is of order $n^{2}$. However, all these FLLN results are about epidemic models on random graphs that are Markovian, where the infectious durations or recovery times are exponentially distributed. In the present paper, we establish an FLLN for a non-Markovian SIR model on large random graphs that results in a measure-valued limit on a graphon.
@@
Non-Markovian stochastic epidemic models with a homogeneous population (no age structure, homogeneous social connectivity, etc.) have been recently studied, see the survey [26]. Although these models offer much more possibilities to fit observational data on infection profiles, many of the classical tools in probability cannot be directly exploited. In particular, the standard epidemic models with a constant infection rate and a general infection duration distribution are recently studied in [41] (see also [53, 52] for Gaussian approximations and [47] for measure-valued state descriptors and PDE limit for the FLLN), and epidemic models with varying infectivity are studied in [24, 42, 25], where deterministic or stochastic integral equations are derived for the FLLNs and FCLTs, respectively. When adding the age of infection into the variables of the model, measure-valued processes have been used to describe the epidemic evolution dynamics in models with such infection-age dependent infectivity, and the corresponding PDE and SPDE limits are established for the FLLNs and FCLTs in [43] and [44], respectively. An epidemic model with contact tracing and general infection duration distributions is studied in [20], where measure-valued processes are used to describe the dynamics and the FLLN is established with a PDE limit. In [37], an individual-based multi-layer SIR model with households and workplaces to take into account the social connectivity heterogeneity is studied, where measure-valued processes tracking remaining infectious times are used to describe the epidemic dynamics and a PDE limit is established for the FLLN.
@@
In this paper we consider an SIR model with infection-age dependent infectivity on a large random graph that captures both individual and connectivity heterogeneities. In particular, connectivities among individuals are given by a non-homogeneous random graph. Each node on this graph represents a single individual, each associated with an individual characteristic/feature. These may account for spatial locations, age and/or social behaviors. So they may be distributed on a compact subset of ${\mathbb{R}}^{d}$, as in the spatial SIR model considered in [34], which we generalize in the simplest setting of pairwise interactions. In the case of age structure or social activity levels, our model generalizes [11] in that they only consider a finite number of compartments while our model includes a possibly continuous age model. The random connectivity of each edge is then assumed to depend on the individuals’ characteristics in a general manner. The connectivities indicate various levels of social activities, which may for instance depend on the spatial locations or age. The random graphs that we consider are of the same form as the ones in [18]. The main difference is that the connectivity of each edge was assumed in [18] to be a deterministic function of the pair of individuals’ characteristics, while we allow for an additional degree of randomness. In [35], there was no heterogeneity in the contact rate and the dependency of the graph on individual types is more restrictive. Our model is complementary to the framework of stochastic block models considered in [10].
@@
Each individual is moreover associated with a random varying infectivity function, which reflects the force of infection of each individual at any time after infection, and depends as well on the individual characteristics. As a result, the infectious duration is also individualized (we express the corresponding distribution as a function $F_{x}(\cdot)$ of the individual characteristic $x$).
@@
Individuals are grouped into three compartments: Susceptible, Infected and Recovered. Infections are generated by the interactions between susceptible and infected individuals. Each individual is thus exposed to a specific force of infection which changes over time and is given by aggregating the weighted infectivity levels of the individuals that are connected to him/her in the graph. The whole dynamic of the epidemic depends on the non-homogeneous random graph and on the random infectivity functions through the expression of this total force of infection for each individual.
@@
To describe the evolution dynamics, we use three measure-valued processes for the three compartments, where at each time $t$ the measure for the infected process is over both the individual characteristics and the infection ages, while the measure for the susceptible and recovered processes is only over the individual characteristics. We prove a FLLN for these measure-valued processes when the population size goes to infinity. As the population size increases, so does the connectivity graph. We impose conditions on the connectivity probabilities to keep the graph consistency as it grows, which notably covers the case of a graphon in the limit. The limits for the FLLN are given by a set of measure-valued equations, from which we further derive a PDE model for the measure-valued infection process. The PDE model is linear, but with a boundary condition which is given by the product of the measure-valued susceptible process and the aggregate force of infection. To note, the dynamics described from the set of measure-valued equations and the PDE may be seen as evolving on a graphon.
@@
The proof for the FLLN is challenging with this level of realism because of the complicated dependence among the variables and processes. First of all, the individual characteristics are not assumed to be i.i.d., and they affect the random infectivity functions and the edge connectivities. Second, the measure-valued processes are driven by the aggregate forces of infection with which infectious individuals transmit the disease to susceptible ones, thereby propagating the epidemic. Notably, these forces of infection are not functions of the three measure-valued state descriptors only, but depend also upon the detailed graph structure. Thus it is difficult to prove the convergence of the processes directly.
@@
Our approach consists in constructing an intermediate model for which the infection rates are derived from the limiting model of the FLLN rather than the interactions through the connectivity graph. This approach facilitates the proof of convergence by enabling specific decoupling of the dependencies. It extends the one that was first developed for the homogeneous model in [25], which was motivated by ideas from the theory of propagation of chaos [50]. In this way, the regularity conditions imposed on the random varying infectivity functions in [24] could be relaxed to only require their boundedness. The current extension is highly non-trivial however because of the heterogeneities among both the individuals and the connectivities. Moreover, we consider measure-valued processes in this paper, while in [25], only total count processes are considered.
@@
In addition, we have made the effort to allow for the individual characteristics to take values in a state space as general as possible instead of the classical choice of $[0,1]$ for graphon kernels. This is particularly of importance regarding the regularity assumptions of the interaction kernel in the limit, which we only assume to be almost everywhere continuous. One of our motivations was to allow for kernels on multidimensional state spaces that possibly display discontinuity thresholds in terms of the distance between the two parameters, as discussed in [18]. We prove in Appendix A a generic convergence result of independent interest, to deal with these weak regularity properties in conjunction with weak convergence of random measures in probability. Our approach demonstrates its strength by its ability to treat the level of graph connectivity heterogeneity assumed in the present paper, in addition to the randomness of the infectivity functions. The conditions that we assume on the average infection rate per edge and on the average variance in this infection rate seem very close to optimal. We refer the reader to Section 5 for the detailed construction of the intermediate processes and for the proofs of weak convergence.
@@ 1.1. Organization of the paper
The rest of the paper is organized as follows. We give a summary of some common notations used throughout the paper in the next subsection. We provide a detailed model description and our assumptions in Section 2. In Section 3, we state the main result of the paper. We establish the uniqueness of the solution to the limiting deterministic PDE model and establish some properties of that solution in Section 4. The proofs of the convergence in the FLLN are given in Section 5, with additional technical supporting results proved in Appendix A.
@@ 1.2. Notation
For a measurable set $(\mathcal{X},\mathcal{F})$, we denote by $\mathcal{M}_{1}(\mathcal{X})$ and $\mathcal{M}(\mathcal{X})$ the sets of respectively probability and finite measures on $\mathcal{X}$. For $\mu\in\mathcal{M}(\mathcal{X})$ and a real-valued measurable function $f$ defined on $\mathcal{X}$, we will sometimes denote the integral of $f$ with respect to the measure $\mu$, if well-defined, by $\langle\mu,\,f\rangle=\int_{\mathcal{X}}f(x)\,\mu(\mathrm{d}x)=\int f\,\mathrm{d}\mu$. For a metric space $\mathcal{X}$ endowed with its Borel $\sigma$-field, we endow $\mathcal{M}_{1}(\mathcal{X})$ and $\mathcal{M}(\mathcal{X})$ with the topology of weak convergence. Provided that $\mathcal{X}$ is a Polish metric space (complete and separable), then so is $\mathcal{M}_{1}(\mathcal{X})$ and $\mathcal{M}(\mathcal{X})$ with the Lévy-Prokhorov metric, cf. [45, Section II.4]. We also denote by $\mathcal{D}(\mathbb{R}_{+},\mathcal{X})$ the space of paths from $\mathbb{R}_{+}$ to $\mathcal{X}$ that are right-continuous with left limits (càd-làg). This space is endowed with the Skorokhod topology (see, e.g., [8, Chapter 3]). In particular, for the measure-valued processes, we denote by $\mathcal{D}(\mathbb{R}_{+},\mathcal{M}(\mathcal{X}))$ the space of càd-làg paths from $\mathbb{R}_{+}$ to $\mathcal{M}(\mathcal{X})$.
In what follows $(\mathbf{\mathbb{X}},\mathrm{d})$ will denote a Polish metric space.
@@ 2. Model description
We consider a population consisting of $N$ individuals, who are in one of the three states – susceptible, infected or recovered at each time, and may interact with each other according to a random graph described below. Susceptible individuals can be infected randomly through interactions with the infected ones. Once infected, an individual will become infectious for a random period of time until recovery, and once recovered, he or she will no longer infect any susceptibles, nor become infected again. Let $\mathcal{S}^{N}(t),\mathcal{I}^{N}(t),\mathcal{R}^{N}(t)$ be the subsets of [[1,N]][\![1,N]\!] that denote the sets of susceptible, infected or recovered individuals at time $t$, respectively. The corresponding processes $S^{N}(t)=|\mathcal{S}^{N}(t)|$, $I^{N}(t)=|\mathcal{I}^{N}(t)|$ and $R^{N}(t)=|\mathcal{R}^{N}(t)|$ denote the numbers of susceptible, infected or recovered individuals at time $t$.
@@ 2.1. Infection-age dependent infectivity
For an individual $i\in\mathcal{S}^{N}(0)$, let $\tau_{i}^{N}>0$ be the time at which he/she starts being infected, and let $A_{i}^{N}(t)=t-\tau_{i}^{N}$ be the infection age of the individual $i$ at time $t$ (by default it is equal to zero for $t<\tau_{i}^{N}$). For an initially infected individual $j\in\mathcal{I}^{N}(0)$, let $\tau_{j}^{N}=-A_{j}^{N}(0)$ denote the infection time before time 0, so that $A_{j}^{N}(t)=t+A_{j}^{N}(0)$ is the infection age at time $t$.
@@
At time $t$, each individual $i\in\mathcal{I}^{N}(t)$ who is infected has an infectivity function depending randomly on its infection age:
$\lambda^{N}_{i}(A_{i}^{N}(t)),\quad t\geq 0\,$, (2.1)
where $\lambda^{N}_{i}$ is assumed not to be identically 0. In particular, for the initially infected individuals $j\in\mathcal{I}^{N}(0)$, their infectivity level at a given time $t>0$ is:
@@
A typical way to define these random functions $\lambda^{N}_{i}(\cdot)$ is specified as follows, although we do not rely in our proofs on this representation. Let $\mathbf{\mathbb{Y}}$ be a Polish space. Let $\mathcal{Y}^{N}=\{Y_{i}^{N}\}_{i\leq N}$ be an i.i.d. sequence of random variables taking values in $\mathbf{\mathbb{Y}}$, with common distribution $\mu_{Y}$. We assume that the sequence $\mathcal{Y}^{N}=\{Y_{i}^{N}\}_{i\leq N}$ is independent from $\mathcal{F}_{0}^{N}$. Let $\widehat{\lambda}$ be a deterministic measurable function from $\mathbf{\mathbb{X}}\times\mathbf{\mathbb{Y}}\times\mathbb{R}_{+}$ to $\mathbb{R}_{+}$ that is càd-làg in its last component. Then we can define the random functions $\lambda^{N}_{i}(\cdot)$ for each $i$ as
@@
$\eta^{N,0}_{j}:=\sup\{a>0:\lambda_{j}^{N}(A^{N}(0)+a)>0\}\,$,
while assuming that there is no pathological situation where the above set would be empty. We specify the individual distribution of infection duration through its cumulative distribution function. Under the formalism of Remark 2.1 it is to be thought as depending only on the variables $(X_{i}^{N},Y^{N}_{i})$.
@@
There exists a measurable function $(F_{x}(a))_{x\in\mathbf{\mathbb{X}},a\in\mathbb{R}_{+}}$ with values in $[0,1]$ which is right-continuous, non-decreasing in the variable ‘$a$’ such that $\lim_{a\to\infty}F_{x}(a)=1$ for all $x\in\mathbf{\mathbb{X}}$ and such that the two following identities hold a.s. for any $N\geq 1$, $t>0$, $i\in\mathcal{S}^{N}(0)$ and $j\in\mathcal{I}^{N}(0)$ with the notation $F^{c}_{x}(a)=1-F_{x}(a)$:
@@
Recall that $\mathbb{P}_{0}^{N}(\eta^{N}_{i}>t)$ equals by definition $\mathbb{P}\big(\eta^{N}_{i}>t\,\big|\,\mathcal{F}_{0}^{N}\big)$ and similarly that $\mathbb{P}_{0}^{N}(\eta^{N,0}_{j}>t)=\mathbb{P}\big(\eta^{N,0}_{j}>t\,\big|\,\mathcal{F}_{0}^{N}\big)$. Note that our previous assumptions on $\eta^{N,N}$ prevents the denominator in the above expression from being 0.
@@
This assumption appears more explicitly stated under the formalism presented in Remark 2.1. By definition, there exists a deterministic measurable function $g:\mathbf{\mathbb{X}}\times\mathbf{\mathbb{Y}}\to\mathbb{R}_{+}$ such that $\eta^{N}_{i}=g(X^{N}_{i},Y^{N}_{i})$ a.s. For each $x\in\mathbf{\mathbb{X}}$, we define the cumulative distribution:
@@
For any $N$ and any initially infected individual $j\in\mathcal{I}^{N}(0)$, the constraint that $\eta^{N,0}_{j}$ is a well-defined positive quantity, which translates into $g(X_{j}^{N},Y^{N}_{j})>A_{j}^{N}(0)$, explains the form of the second identity in Assumption 2.4, where for any $t>0$,
@@
There exists a deterministic measurable function $\bar{\lambda}$ from $\mathbf{\mathbb{X}}\times\mathbb{R}_{+}$ to $\mathbb{R}_{+}$ such that the two following identities hold for any $N\geq 1$, any $a\geq 0$, any $t\geq 0$, any $i\in\mathcal{S}^{N}(0)$ and any $j\in\mathcal{I}^{N}(0)$: